I need Your Help in drawing a pie chart ?
please help and tell me how to make a pie chart.
I Shall be very thankful to you because i am taking a exam of maths tomorrow.

Answer by stanbon(75887)

I need Your Help in drawing a pie chart ?
please help and tell me how to make a pie chart.
-------------------
Let's say you have:
A: 12 items
B: 17 items
C: 25 items
---------------
1st: Add the items to get 54
----------
2nd: Convert each item count to percentage:
12/54 = 22%
17/54 = 31%
25/54 = 46%
----------------
3rd: Draw a circle and estimate the percentages as pieces of pie
that fill the circle.
===============================
Cheers,
Stan H.